# Do Vision-Language Models Represent Space and How? Evaluating Spatial Frame of Reference under Ambiguities

### v1

@@ 1 Introduction
The recent success of large language models has sparked breakthroughs in multi-modalities, leading to the development of many vision-language models (VLMs; Chen et al., 2023b; OpenAI, 2024; Reid et al., 2024, inter alia). With some benchmarks developed to evaluate the downstream performance of these models (Liu et al., 2023c; Yue et al., 2024), there has been growing excitement around evaluations and analyses inspired by human cognitive capabilities such as referential grounding (Ma et al., 2023a), compositional reasoning (Ma et al., 2023c), visual illusions (Zhang et al., 2023; Guan et al., 2024), and theory of mind (Jin et al., 2024). One direction among them that captures significant attention is spatial language understanding and reasoning, leading to several benchmarks (Kamath et al., 2023; Liu et al., 2023a) and enhanced models (Chen et al., 2024a; Cheng et al., 2024).
@@ 2.1 Spatial Language and Spatial Representation
Ambiguities in relative FoRs. The variations of relative FoRs form another source of ambiguity. After putting the origin of the coordination system on the viewer, multiple strategies specifying how to transform the axes can be considered (Figure 1b). Different languages use different transformation conventions to resolve the ambiguity of the front-back and left-right of a relatum (Levinson, 2003; Shusterman & Li, 2016), including: (1) translated projection (e.g., Hausa) where the coordinate frame of the speaker is directly applied, (2) rotated projection (e.g., Tamil), where the coordinate frame of the speaker is transformed with a 180-degree rotation, and (3) reflected projection (e.g., English), where only the front-back axis is reversed.
2.2 Spatial Understanding in Vision-Language Models
@@ 3.4 Metrics
Quantitatively assessing the spatial understanding and reasoning capabilities of models is challenging for two reasons. First, the physical world is continuous, and spatial relations may extend beyond the precise canonical front-back and left-right axes. As noted by Carlson-Radvansky & Logan (1997), there exists regions of acceptability where, for instance, an object slightly to the front-left might still be considered being “in front.” Second, as Dentella et al. (2023) pointed out, language models are biased towards affirmative responses. However, the intermediate representations may be sensitive to variations in input and, to some extent, align with human perceptions of spatial cues. Based on these concerns and findings, we introduce multiple metrics to evaluate the models’ competence to enable more nuanced analyses in addition to accuracy that measures performance.
@@
Accuracy. Given a spatial relation in the textual prompt, we assess whether the assigned response probabilities correspond to whether the referent lies within the acceptable region defined by the relatum and . Formally, we define the local probability of the model responding ‘Yes’ by . We consider the inference correct if (1) the scene falls into the acceptability region and or (2) the scene falls out of the acceptability region and .
@@ 4.2 Most VLMs Prefer Egocentric Relative Frame of Reference
We now attempt to answer the research question: do VLMs have a preferred frame of reference, and if so, what is it? We conduct our study on COMFORT-CAR using the no perspective prompt (nop) that deliberately leaves the FoR ambiguous. When calculating the performance with respect to relative FoRs (either egocentric or addressee-centered), we assume a reflected coordinate transformation convention. Table 3 shows the results of preferred FoR in English measured by the region parsing error , and Table 8 in the appendix shows the complete evaluation including both and . Almost all VLMs demonstrate a significant preference for the egocentric relative FoR similar to English, again, except for the BLIP series. Additionally, the models’ performances are inconsistent across different spatial relations—models generally perform better in the lateral directions (left and right) than the sagittal ones (front and behind), even in competitive industry models like GPT-4o. For instance, GLaMM does not show a very strong preference when resolving ambiguities along the sagittal axes, but it demonstrates a significant preference when resolving lateral ambiguity.
@@ 4.5 A Cross-lingual and Cross-cultural Evaluation of Frame of Reference
All previous experiments are centered around English; however, individuals from diverse multilingual and cultural backgrounds may adopt different preferences and conventions to select their FoR in resolving ambiguities (Majid et al., 2004; O’Meara & Báez, 2011; Bohnemeyer et al., 2014; Ogelo & Bylund, 2024). Our next research question naturally arises: Does multilingual VLMs faithfully follow the preferences and conventions (associated with different languages) to select the FoR? To extend the study of preferred FoR from English to a multilingual setting, we evaluate 109 languages worldwide to investigate whether each language shows a preferred FoR. We translate the English prompts into the target languages using the Google Cloud Translate API. Given that the open-source language models either lack strong multilingual capabilities or underperform in previous evaluations, we study this problem on the GPT-4o model (OpenAI, 2024). We follow the setup similar to Section 4.2, but only evaluate the images corresponding to the four canonical directions using the nop prompt. For each language, we compute for each FoR and coordinate transformation. Figure 8 presents a visualization of the world map, displaying the preference of each region for using the (object-centered) intrinsic FoR over the relative FoR, where the latter corresponds to a low value. Table 10 summarizes the results for all languages tested.
@@ 5.1 Do vision-language models represent space and how?
It is insufficient to answer this question by simply querying the model with text-image pairs and comparing the output with a fixed ground truth. We must, at least, query the models with awareness of the ambiguity in FoRs, which is essential in determining how the scenes in the physical world are mapped to spatial expressions (Levinson, 2003). Our experiments confirm that many VLMs are equipped with reasonable spatial representations through vision-language training alone; in particular, most VLMs clearly prefer the egocentric relative FoR with reflected projection, aligning with English conventions. However, our results also show these representations lack robustness and consistency in a continuous space. Similar experimental setups can yield widely varying performance across different spatial relations—for example, GPT-4 shows minimal preferences for the egocentric relative FoR along the sagittal axis but a significant preference along the lateral one (Table 3). As a result, VLMs demonstrate unsatisfactory consistency in their spatial performance (Table 5). Future work is necessary to improve the consistency and robustness of spatial representations in these models.
@@ 5.2 Perspective taking as a prerequisite of human-like spatial reasoning
Most languages support multiple FoRs.333Some languages, in very rare cases, have only one available spatial frame of reference. For example, Guugu Yimithirr exclusively uses the absolute FoR (Levinson, 2003). The ability to understand and reason about space from a non-egocentric perspective is an important foundation of the Theory of Mind, a basic building block of our situated communication skill that allows us to infer others’ mental states (Ma et al., 2023b). One of our key findings is that VLMs still struggle to adopt alternative FoRs flexibly, even when provided with explicit perspective-taking instructions (Section 4.3). We hypothesize that this phenomenon may come from a reporting bias in the image-text datasets available on the internet—it is natural to take the reflected relative FoR to view images presented on a screen, but this does not always apply in real-world applications. To address this issue, we suggest future work extend the current 2D VLMs to the 3D domain, by considering camera poses and multiview data (Yang et al., 2024) for training.
@@ 5.3 Cross-cultural conventions in cross-lingual spatial understanding
The conventions for resolving spatial ambiguities are not uniform, as individuals from diverse linguistic and cultural backgrounds select their FoR differently. Cultural conventions can even be transmitted as individuals are exposed to other languages. Interestingly, Bohnemeyer et al. (2014) found that among native speakers of Indigenous languages (with various preferences in FoRs), those more proficient in Spanish tend to use the reflected relative FoR (Spanish convention) more in their native language. This phenomenon has led to their Linguistic Transmission Hypothesis: “Using any language or linguistic variety – independently of its structures – may facilitate the acquisition of cultural practices of nonlinguistic cognition shared among the speakers of the language.” Analogously, our experiment raises important concerns that English may dominate the FoR preference conventions of other languages in multilingual VLMs. This is not surprising, as current training recipes for multilingual multimodal language models heavily rely on machine-translated captions (Chen et al., 2023b; Geigle et al., 2024). However, this practice can be problematic. For instance, Hausa prefers an interpretation where the “front” aligns with the English concept of “back,” (Hill, 1982), where this approach may lead to English conventions overshadowing those of other languages. At a high level, this issue is not limited to spatial reasoning—for example, Shi et al. (2023) have demonstrated that English is always the best chain-of-thought language for math reasoning with multilingual LLMs, no matter what language is used for the problem description. To enable similar linguistic transmission in AI models, exposure to naturally generated multilingual data is crucial (Romero et al., 2024).
@@ Acknowledgments
This work was supported in part by NSF IIS-1949634, NSERC RGPIN-2024-04395, and the Microsoft Accelerate Foundation Models Research (AFMR) grant program. Ziqiao Ma is supported in part by the Weinberg Cognitive Science Fellowship. The authors would like to thank Yinpei Dai, Run Peng, Jung-Chun Liu, and Xuejun Zhang for proofreading and feedback.
@@ Appendix B Addendum to Results
| kn | Kannada | 52.3 | 40.8 | 53.2 | 49.5 | 49.6 | 51.9 | 53.4 | zh | Yoruba | 54.6 | 35.5 | 58.3 | 51.4 | 56.9 | 57.2 | 57.8 |
| ko | Korean | 53.6 | 36.5 | 59.1 | 49.7 | 53.3 | 53.5 | 59.8 | zu | Chinese | 55.6 | 35.9 | 57.7 | 53.1 | 55.3 | 56.5 | 60.2 |

### v2

@@ 1 Introduction
The recent success of large language models has sparked breakthroughs in multi-modalities, leading to the development of many vision-language models (VLMs; Chen et al., 2023b; OpenAI, 2024; Reid et al., 2024, inter alia). With some benchmarks developed to evaluate the downstream performance of these models (Liu et al., 2023c; Yue et al., 2024), there has been growing excitement around evaluations and analyses inspired by human cognitive capabilities such as referential grounding (Ma et al., 2023a), compositional reasoning (Ma et al., 2023c), visual illusions (Zhang et al., 2023; Guan et al., 2024), and theory of mind (Jin et al., 2024). One direction among them that captures significant attention is spatial language understanding and reasoning, leading to several benchmarks (Mirzaee et al., 2021; Mirzaee & Kordjamshidi, 2022; Kamath et al., 2023; Liu et al., 2023a) and enhanced models (Chen et al., 2024a; Cheng et al., 2024; Premsri & Kordjamshidi, 2025).
@@ 2.1 Spatial Language and Spatial Representation
Ambiguities in relative FoRs. The variations of relative FoRs form another source of ambiguity. After putting the origin of the coordination system on the viewer, multiple strategies specifying how to transform the axes can be considered (Figure 1b). Different languages adopt different transformation conventions to relative FoRs (Levinson, 2003; Shusterman & Li, 2016), including: (1) translated projection (e.g., Hausa) where the coordinate frame of the speaker is directly applied, (2) rotated projection (e.g., Tamil), where the coordinate frame of the speaker is transformed with a 180-degree rotation, and (3) reflected projection (e.g., English), where only the front-back axis is reversed.
2.2 Spatial Understanding in Vision-Language Models
@@ 3.4 Metrics
Quantitatively assessing the spatial understanding and reasoning capabilities of models is challenging for two reasons. First, the physical world is continuous, and spatial relations may extend beyond the precise canonical sagittal and lateral axes. As noted by Carlson-Radvansky & Logan (1997), there exists regions of acceptability where, for instance, an object slightly to the front-left might still be considered being “in front.” Second, language models are biased towards affirmative responses (Dentella et al., 2023). However, the intermediate representations may be sensitive to variations in input and, to some extent, align with human perceptions of spatial cues. Based on these concerns and findings, we introduce multiple metrics to evaluate the models’ competence to enable more nuanced analyses in addition to accuracy that measures performance.
@@
Accuracy. Given a spatial relation in the textual prompt, we assess whether the assigned response probabilities correspond to whether the referent lies within the acceptable region defined by the relatum and . In line with recent work (Hu & Levy, 2023; Wang & Shi, 2025), we define the local probability of the model responding ‘Yes’ by . We consider the inference correct if (1) the scene falls into the acceptability region and or (2) the scene falls out of the acceptability region and .
@@ 4.2 Most VLMs Prefer Egocentric Relative Frame of Reference
We now attempt to answer: do VLMs have a preferred frame of reference, and if so, what is it? We conduct our study on COMFORT-CAR using the no perspective prompt (nop) that deliberately leaves the FoR ambiguous. When calculating the performance with respect to relative FoRs (either egocentric or addressee-centered), we assume a reflected coordinate transformation convention. Table 3 shows the results of preferred FoR in English measured by the region parsing error , and Table 8 in the appendix shows the complete evaluation including both and . Almost all VLMs demonstrate a significant preference towards the egocentric relative FoR similar to English, again, except for the BLIP series. Additionally, the models’ performances are inconsistent across different spatial relations—models generally perform better in the lateral directions (left and right) than the sagittal ones (front and behind), even in competitive industry models like GPT-4o. For instance, GLaMM does not show a very strong preference when resolving ambiguities along the sagittal axes, but it demonstrates a significant preference when resolving lateral ambiguity.
@@ 4.5 A Cross-lingual and Cross-cultural Evaluation of Frame of Reference
All previous experiments are centered around English; however, individuals from diverse multilingual and cultural backgrounds may adopt different preferences and conventions to select their FoR in resolving ambiguities (Majid et al., 2004; O’Meara & Báez, 2011; Bohnemeyer et al., 2014; Ogelo & Bylund, 2024). Our next research question naturally arises: Do multilingual VLMs faithfully follow the preferences and conventions (associated with different languages) to select the FoR? To extend the study of preferred FoR from English to a multilingual setting, we evaluate 109 languages worldwide to investigate whether each language shows a preferred FoR. We translate the English prompts into the target languages using the Google Cloud Translate API. Given that the open-source language models either lack strong multilingual capabilities or underperform in previous evaluations, we study this problem on the GPT-4o model (OpenAI, 2024). We follow the setup similar to Section 4.2, but only evaluate the images corresponding to the four canonical directions using the nop prompt. For each language, we compute for each FoR and coordinate transformation. Figure 8 presents a visualization of the world map, displaying the preference of each region for using the (object-centered) intrinsic FoR over the relative FoR, where the latter corresponds to a low value. Table 10 summarizes the results for all languages tested.
@@ 5.1 Do vision-language models represent space and how?
It is insufficient to answer this question by simply querying the model with text-image pairs and comparing the output with a fixed ground truth. We must, at least, query the models with awareness of the ambiguity in FoRs, which is essential in determining how the scenes in the physical world are mapped to spatial expressions (Levinson, 2003). Our experiments confirm that many VLMs are equipped with reasonable spatial representations through vision-language training alone; in particular, most VLMs clearly prefer the egocentric relative FoR with reflected projection, aligning with English conventions. However, our results also show these representations lack robustness and consistency in a continuous space. Similar experimental setups can yield widely varying performance across different spatial relations—for example, GPT-4o shows minimal preferences for the egocentric relative FoR along the sagittal axis but a significant preference along the lateral one (Table 3). As a result, VLMs demonstrate unsatisfactory consistency in their spatial performance (Table 5). Future work is necessary to improve the consistency and robustness of spatial representations in these models.
@@ 5.2 Perspective taking as a prerequisite of human-like spatial reasoning
Most languages support multiple FoRs.333Some languages, in very rare cases, have only one available spatial frame of reference. For example, Guugu Yimithirr exclusively uses the absolute FoR (Levinson, 2003). The ability to understand and reason about space from a non-egocentric perspective is an important foundation of the Theory of Mind, a basic building block of our situated communication skill that allows us to infer others’ mental states (Ma et al., 2023b). One of our key findings is that VLMs still struggle to adopt alternative FoRs flexibly, even when provided with explicit perspective-taking instructions (Section 4.3). We hypothesize that this phenomenon may come from a reporting bias in the image-text datasets available on the internet—it is natural to take the reflected relative FoR to view images presented on a screen, but this does not always apply in real-world applications. To address this issue, we suggest future work extend the current 2D VLMs to the 3D domain, by considering camera poses and multiview data (Yang et al., 2024; Hao et al., 2024) for training.
@@ 5.3 Cross-cultural conventions in cross-lingual spatial understanding
The conventions for resolving spatial ambiguities are not uniform, as individuals from diverse linguistic and cultural backgrounds select their FoR differently. Cultural conventions can even be transmitted as individuals are exposed to other languages. Interestingly, Bohnemeyer et al. (2014) found that among native speakers of Indigenous languages (with various preferences in FoRs), those more proficient in Spanish tend to use the reflected relative FoR (Spanish convention) more in their native language. This phenomenon has led to their Linguistic Transmission Hypothesis: “Using any language or linguistic variety, independently of its structures, may facilitate the acquisition of cultural practices of nonlinguistic cognition shared among the speakers of the language.” Analogously, our experiment raises important concerns that English may dominate the FoR preference conventions of other languages in multilingual VLMs. This is not surprising, as current training recipes for multilingual multimodal language models heavily rely on machine-translated captions (Chen et al., 2023b; Geigle et al., 2024). However, this practice can be problematic. For instance, Hausa prefers an interpretation where the “front” aligns with the English concept of “back,” (Hill, 1982), where this approach may lead to English conventions overshadowing those of other languages. At a high level, this issue is not limited to spatial reasoning—for example, Shi et al. (2023) have demonstrated that English is always the best chain-of-thought language for math reasoning with multilingual LLMs, no matter what language is used for the problem description. To enable similar linguistic transmission in AI models, exposure to naturally generated multilingual data is crucial (Romero et al., 2024; Nguyen et al., 2024).
@@ Acknowledgments
This work was supported in part by NSF IIS-1949634, NSF SES-2128623, NSERC RGPIN-2024-04395, ONR N00014-23-1-2417, and the Microsoft Accelerate Foundation Models Research (AFMR) grant program. Ziqiao Ma is supported in part by the Weinberg Cognitive Science Fellowship. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of these funding agencies. The authors would like to thank Yinpei Dai, Run Peng, Jung-Chun Liu, and Xuejun Zhang for proofreading and feedback.
@@ Appendix B Addendum to Results
| kn | Kannada | 52.3 | 40.8 | 53.2 | 49.5 | 49.6 | 51.9 | 53.4 | zh | Chinese | 54.6 | 35.5 | 58.3 | 51.4 | 56.9 | 57.2 | 57.8 |
| ko | Korean | 53.6 | 36.5 | 59.1 | 49.7 | 53.3 | 53.5 | 59.8 | zu | Zulu | 55.6 | 35.9 | 57.7 | 53.1 | 55.3 | 56.5 | 60.2 |